OpenStudy (anonymous):

Divide 4 - 5i/1 + 4i Answer is -16/17,-21/17i

OpenStudy (anonymous):

hmmm there is only one answer multiply by 1-4i numerator and denominator the numerator will become (4-5i)(1-4i) = 4-16i-5i-20 = -21i-16 and the denominator will become (1+4i)(1-4i) = 17 so the answer is (-21i-16)/17

OpenStudy (anonymous):

multiply by 1-4i both numerator and denominator *